OpenStudy (anonymous):

Arlene is showing her work in simplifying −7.9 + 8.2 − 3.4 + 2.1. Identify any errors in her work or in her reasoning. Write feedback to Arlene explaining the mistake she made and how to correct her work.

OpenStudy (anonymous):

what did she solve

OpenStudy (anonymous):

i only see the question not what she did =)

OpenStudy (anonymous):

Step Work Reasoning −7.9 + 8.2 − 3.4 + 2.1 Original problem 1 −7.9 + 8.2 + (−3.4) + 2.1 Additive inverse 2 −7.9 + (−3.4) + 8.2 + 2.1 Associative property 3 (−7.9 + (−3.4)) + (8.2 + 2.1) Commutative property 4 −11.3 + 10.3 Simplify 5 −1 Simplify

OpenStudy (anonymous):

......

OpenStudy (anonymous):

=)

OpenStudy (anonymous):

Keep doing each step manually, and see where it slips up: −7.9+8.2−3.4+2.1=−1 −7.9+8.2+(−3.4)+2.1=−1 −7.9+(−3.4)+8.2+2.1=−1 (−7.9+(−3.4))+(8.2+2.1)=−1 −11.3+10.3=−1

OpenStudy (anonymous):

bro this one

OpenStudy (anonymous):

Pedro deposited $35 in his savings account, then withdrew $50.

OpenStudy (anonymous):

what about it ?

OpenStudy (anonymous):

he cant because then he would owe the bank =)

OpenStudy (anonymous):

its need a number line

OpenStudy (anonymous):

so start with zero go up 35 then subtract 50 so go down to -15

OpenStudy (anonymous):

ok bro

OpenStudy (anonymous):